# SuperAD: A Training-free Anomaly Classification and Segmentation Method for CVPR 2025 VAND 3.0 Workshop Challenge Track 1: Adapt & Detect

### v1

@@ 2.2.1 Dataset Utilization
Can 58.40 0.15
Fabric 69.06 28.22
Fruit Jelly 79.71 49.89
Rice 92.44 66.59
Vial 69.46 36.81
Wallplugs 77.95 19.82
Walnuts 90.03 75.65
Sheet Metal 76.84 39.76
Mean 76.74 39.61
@@
Can 0.3 / 0.1 0.1 / 0.1 0.1 / 0.1 0.8 / 0.1 5.0 / 0.1 0.6 / 0.1 0.4 / 0.1 17.9 / 1.8
Fabric 11.5 / 9.8 2.6 / 2.2 2.9 / 2.3 7.6 / 1.0 22.0 / 4.1 21.6 / 10.2 7.9 / 5.0 77.6 / 59.7
Fruit Jelly 8.7 / 8.2 22.5 / 22.7 26.9 / 26.7 20.8 / 18.2 47.6 / 38.1 25.1 / 23.0 17.9 / 17.2 45.1 / 44.5
Rice 3.8 / 4.2 7.0 / 3.9 9.5 / 2.9 15.0 / 0.5 19.1 / 1.8 11.6 / 1.0 1.5 / 1.4 60.1 / 60.3
Sheet Metal 1.8 / 1.1 41.3 / 39.2 40.9 / 37.7 9.3 / 3.8 13.0 / 7.6 14.6 / 2.8 13.9 / 14.4 58.8 / 58.9
Vial 2.3 / 2.2 28.0 / 28.3 28.2 / 22.8 30.5 / 26.5 23.3 / 6.2 31.9 / 17.5 28.2 / 27.9 41.5 / 39.7
Wallplugs 0.0 / 0.0 1.9 / 0.8 1.3 / 0.9 4.4 / 0.3 0.1 / 0.2 1.0 / 0.3 0.4 / 0.4 7.5 / 6.3
Walnuts 1.2 / 1.3 41.2 / 36.7 44.1 / 40.5 34.6 / 13.3 44.5 / 14.3 35.2 / 14.3 17.0 / 9.6 69.0 / 68.9
Mean 3.7 / 3.4 18.1 / 16.7 19.2 / 16.7 15.4 / 8.0 21.8 / 9.0 17.7 / 8.7 10.9 / 9.5 47.2 / 42.5

### v2

@@ 2.2.1 Dataset Utilization
Can 58.61 0.18
Fabric 68.29 28.22
Fruit Jelly 80.93 48.27
Rice 92.92 68.44
Vial 69.04 35.88
Wallplugs 77.90 19.19
Walnuts 89.23 75.05
Sheet Metal 76.80 40.13
Mean 76.71 39.42
@@
Can 0.3 / 0.1 0.1 / 0.1 0.1 / 0.1 0.8 / 0.1 5.0 / 0.1 0.6 / 0.1 0.4 / 0.1 17.3 / 1.9
Fabric 11.5 / 9.8 2.6 / 2.2 2.9 / 2.3 7.6 / 1.0 22.0 / 4.1 21.6 / 10.2 7.9 / 5.0 77.4 / 65.3
Fruit Jelly 8.7 / 8.2 22.5 / 22.7 26.9 / 26.7 20.8 / 18.2 47.6 / 38.1 25.1 / 23.0 17.9 / 17.2 41.3 / 40.9
Rice 3.8 / 4.2 7.0 / 3.9 9.5 / 2.9 15.0 / 0.5 19.1 / 1.8 11.6 / 1.0 1.5 / 1.4 60.9 / 61.2
Sheet Metal 1.8 / 1.1 41.3 / 39.2 40.9 / 37.7 9.3 / 3.8 13.0 / 7.6 14.6 / 2.8 13.9 / 14.4 59.5 / 59.7
Vial 2.3 / 2.2 28.0 / 28.3 28.2 / 22.8 30.5 / 26.5 23.3 / 6.2 31.9 / 17.5 28.2 / 27.9 42.8 / 40.8
Wallplugs 0.0 / 0.0 1.9 / 0.8 1.3 / 0.9 4.4 / 0.3 0.1 / 0.2 1.0 / 0.3 0.4 / 0.4 13.7 / 6.7
Walnuts 1.2 / 1.3 41.2 / 36.7 44.1 / 40.5 34.6 / 13.3 44.5 / 14.3 35.2 / 14.3 17.0 / 9.6 69.1 / 69.1
Mean 3.7 / 3.4 18.1 / 16.7 19.2 / 16.7 15.4 / 8.0 21.8 / 9.0 17.7 / 8.7 10.9 / 9.5 47.8 / 43.2